Logiweb codex of append

Up Help append, append(∗, ∗),

append

[append →^{pyk} “append”]

append(∗, ∗)

[append(x, y) →^{val} y : : x] [append(x, y) →^{tex} “append(#1., #2.)”] [append(x, y) →^{pyk} “append ∗ to ∗ end append”] The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-02.UTC:19:54:02.237435 = MJD-53523.TAI:19:54:34.237435 = LGT-4624458874237435e-6